Fermi National Laboratory
 

Accelerator Update for July 20, 2007 - July 23, 2007

 

Friday July 20

The day shift began with the Tevatron (TeV) store 5562 in an end-of-store study, with Operations monitoring the Recycler stash (278.6E10), with the Antiproton Source (Pbar) stacking (46.3E10), with the Linac, Booster, and Main Injector (MI) operating normally, with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams, and with MiniBooNE taking beam. NuMI is off until after the summer shutdown.

Operations terminated store 5562 at 8:01 AM.

At 8:37 AM, cryo system technicians reported that a wet engine at TeV sector E4 had failed. They began replacing the engine.

Operations reported at 11:31 AM, that the new Meson MTest experiment had received all of its sign offs and began taking beam at 12:40 PM.

An MI expert began a slipstacking study at 12:41 PM.

Operations reported at 3:58 PM that the Recycler had stashed antiprotons twice during the day shift.

CHL reported at 5:06 PM that cryo system techs had finished the E4 wet engine replacement and that the TeV sector E4 cryo system had cooled down.

MTest stopped using the beam for night at 8:44 PM.

Operations put the TeV through a dry squeeze at 8:59 PM.

Operations put the TeV into shot setup at 9:31 PM.

Operations established store 5565 at 11:44 PM with an initial luminosity of 229.04E30.

Saturday July 21

The midnight shift began with Operations monitoring store 5565 (193.76E30) and the stash (147.4E10) and the stack (14.5E10), and with MiniBooNE taking beam.

At 4:36 AM, Operations reported that a Linac RF station (LRF5) had tripped off twice. It reset both times.

At 4:41 AM, Operations reported I- Source regulation problems. It would not regulate at its normal 750 kV. Operators swapped out the Command and Regulator/Oscillator modules, but it didn’t help. Operators contacted experts.

Operations reported at 6:12 AM that operators switched to the H- Source. Linac resumed sending beam at 6:34 AM.

Operations established beam to MTest at 6:50 AM.

Operations reported at 7:56 AM that the Recycler had stashed three times during the midnight shift.

At 1:47 PM, MTest requested to change from the 120 GeV running mode to the 16 GeV running mode. Operations could not successfully make the mode switch. An operator called in a Switchyard (SY) expert.

At 2:38 PM, the 750 KeV area buncher tripped off. It reset.

The SY expert successfully set MTest to its 16 GeV running mode at 3:06 PM.

Operations disabled MTest for the night at 6:03 PM.

Operations reported at 7:59 PM that the Recycler had stashed five times during the last 12 hours. Accelerator status: store 5565 was at 40.03E30, the stack was at 21.4E10, and the stash was at 311.7E10.

At 8:23 PM, BRF11 tripped off. It reset.

At 8:51 PM, the Linac buncher tripped off. It reset.

Sunday July 22

The midnight shift began with Operations monitoring store 5565 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

Operations terminated store 5565 at 1 AM.

Operations put the TeV into shot setup at 1:12 AM.

At 1:24 AM, Operations reported that shot setup was on hold due to problems with CIA crates at TeV sector D0. Operators found and fixed a crate power supply card with a blown fuse.

At 2:37 AM, BRF16 tripped off. It reset.

At 2:44 AM, Operations reported that 23E10 of antiprotons were lost from the Recycler due to a problem with the new vertical dampers. The emittances were very large. Experts investigated and confirmed that the problem was with the dampers. The Pelletron tripped off at the same time.

Operations put the TeV through a dry squeeze at 4:39 AM.

Operations put the TeV into shot setup at 5:13 AM.

At 6:30 AM, a Linac Klystron RF station (KRF1) tripped off. It reset.

At 7:18 AM, MIRF15 tripped off. It reset.

Operations established store 5567 at 7:31 AM with an initial luminosity of 163.02E30. The low luminosity was due to fewer antiprotons and a TeV sector C49 separator spark.

Recycler experts began transfer measurements at 7:41 AM with the damper on and off.

MTest resumed taking beam at 7:47 AM.

Operations reported at 7:58 AM that the Recycler had stashed three times during the last 12 hours. Accelerator status: store 5567 was at 141.77E30, the stack was at 19.7E10, and the stash was at 79.1E10.

Recycler experts reported at 10:05 AM that they had switched back to the old dampers.

At 11:42 AM, the H- Source high voltage tripped off. It tripped off several more times while operators tried to bring back up. It finally reset.

At 1:27 PM, a TeV RF station (TRF8) tripped off.

Operations disabled the MTest beam at 6:01 PM.

Operations reported at 7:57 PM that the Recycler had stashed four times during the last 12 hours. Accelerator status: store 5567 was at 53.94E30, the stack was at 41.4E10, and the stash was at 231.2E10.

A TeV RF expert reported that TRF8 had a LCW leak. He repaired the leak and reported that the station was ready to turn on between stores.

At 9:28 PM, LRF4 tripped off for the first of two times in about 30 minutes. It reset.

Monday July 23

The midnight shift began with Operations monitoring store 5567 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

At 12:50 AM, BRF11 tripped off. It reset.

At 5:51 AM, store 5567 aborted due to TeV crate D9 dropping out of the link. There was no quench. Operators replaced a failed power supply for crate D9.

MTest resumed taking beam at 6:22 AM.

Operations put the TeV through a dry squeeze at 7:14 AM.

Operations put the TeV into shot setup at 7:41.

Operations reported at 8 AM that the Recycler had stashed five times during the last 12 hours.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Preaccelerator experts to switch back to the I- Source this afternoon and possibly allow some Booster studies.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/23/2007    email Fermilab